OpenStudy (anonymous):

What is the largest value of "x" (in radians) less than pi satisfying the equation: cos x = cos (6x)

OpenStudy (anonymous):

0

OpenStudy (anonymous):

is it the largest though?

OpenStudy (anonymous):

it's the largest value of x, not the largest value of solution

OpenStudy (turingtest):

right, I misunderstood

myininaya (myininaya):

i'm getting x=4pi/5

myininaya (myininaya):

\[\cos(x+2 \pi n)=\cos(6x) => x+2\pi n=6x\]

myininaya (myininaya):

\[5x=2\pi n => x=\frac{2 \pi n}{5}\]

myininaya (myininaya):

so if n=4 x<pi

myininaya (myininaya):

oops n=2

OpenStudy (anonymous):

Myininaya is right

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

What's the Python solution to this problem? import sympy. sympy.solve(cos(x) = cos(6x))